GlueX TOF Meeting, November 15, 2016

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, November 15
11:00 am EST
JLab: CEBAF Center, Room F326/327


  1. Announcements
    1. New PMTs have arrived. (Mark)
  2. Review of minutes from the previous meeting
  3. Integral Non-Linearity Calibration of CAEN TDCs (Beni)
    1. Test Results from Paul M.
  4. HV Scan Planning
  5. Lucite Shield (Paul)
  6. PMT bases with Amplifiers
  7. Action Item Recap



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Thomas Britton, Brad Cannon, Paul Eugenio, Mark Ito (chair), Simon Taylor

There is a recording of this meeting on the BlueJeans site.

Announcement: New PMTs Have Arrived

The 30 new tubes are here. They are in Mark's office. We will look into shipping them to FSU so that they can go through the same acceptance tests we used for the tubes used in construction.

Review of minutes from the previous meeting

We went over minutes from the meeting on October 18.

  • Timing Test at Reduced Gain. Sasha will look into doing this at FSU.
  • Run Readiness. There are four dead ADC channels in the TOF right now. Beni Zihlmann is investigating.
  • Hot Checkout is done for the TOF. Beni took care of it.

Integral Non-Linearity Calibration of CAEN TDCs

Beni complete his calibration of the INL correction. Among other places, he reported results in the e-log. Constants have been downloaded into the CAEN TDCs and Paul Mattione repeated his RF test with much better results.

Beni has also been looking at the feasibility of applying a correction to data already taken. The procedure depends on the the constants and setting resident of the CAEN before now. He is trying to do a calibration using old data to assess the situation, but the results are a bit confusing at present. He is consulting with experts in the Electronics Group.

HV Scan Planning

There have been no functional changes in the software since the last meeting. Paul Eugenio reminded us that there might be some incompatibilities with ROOT 6 remaining.

Lucite Shield

Paul walked us through his elog entry on the test fit on the front face of the TOF and his simple run plan. He is looking into how much data would be needed to do an FCAL calibration (to assess effects on that detector).

TOF Time Resolution for Charged Tracks

Brad is working on a study. He will aim for presenting results at the next meeting.

Action Items

  1. Fix the 4 dead ADC channels. (Brad)
  2. Look into timing tests at FSU. (Sasha)
  3. Look into shipping PMTs to FSU for testing. (Mark)
  4. Look into a replacement LeCroy 1440 mainframe for FSU. (Mark)